# Memory Gym: Towards Endless Tasks to Benchmark Memory Capabilities of Agents

### v3

@@ 5 Experimental Analysis
We run empirical experiments on Memory Gym using the just introduced TrXL baseline and one based on GRU, both powered by Proximal Policy Optimization (Schulman et al., 2017). Contrary to our previous study (Pleines et al., 2023), we exclude HELM (History comprEssion via Language Models) (Paischer et al., 2022a) and its successor, HELMv2 (Paischer et al., 2022b). The decision stems from HELM’s suboptimal wall-time efficiency, which is six times more expensive than a GRU agent (Pleines et al., 2023), making hyperparameter optimization impractical for our purposes. As we are strongly limited by our compute budget of 25,000 GPU hours, we rather dedicated the resources towards the development of the endless environments and the TrXL baseline. We also conduct preliminary tests using Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) on the finite environments without any extensive hyperparameter tuning. These models usually underperformed when applied directly, suggesting that out-of-the-box applications might necessitate hyperparameter optimization.
@@
As part of our evaluation protocol, the following sample efficiency curves depict the interquartile mean (IQM) along with a stratified bootstrapped confidence interval of 0.95 (shaded area on the plots), following the guidelines of Agarwal et al. (2021). Each training run is repeated 5 times. To evaluate the generalization ability of our baselines, we extensively test them on 50 novel environment seeds that were excluded from the training process. This evaluation is repeated 3 times for each data point, resulting in a comprehensive assessment based on a total of 750 episodes for every data point presented.
The next subsection presents a two-fold examination of the finite environments. Firstly, we reaffirm that memory is imperative in simpler environment instances, a conclusion previously drawn in our earlier study (Pleines et al., 2023). Enhanced hyperparameters allow us to report improved results. Secondly, we compare the performance of GRU and TrXL on the original finite tasks. TrXL exhibits stronger policies in Mortar Mayhem than GRU, while it is more sample efficient in Mystery Path. Searing Spotlights is an exception to this where TrXL is less sample efficient. Contradicting our initial expectations and the results on the finite environments, we unveil surprising findings that highlight GRU’s superiority over TrXL in all endless environments. For a comprehensive overview of the hyperparameters employed in our experiments, see Appendix B. Results for GTrXL and LSTM on the finite environments can be found in Appendix C. Additionally, a selection of wall-time efficiency statistics is available in Appendix D.
@@ 5.1 Finite Environments: Transformer-XL is often more efficient than GRU
Figure 6: Performance comparison across instances of Memory Gym’s finite environments. Mortar Mayhem Act Grid (a) provides commands as a fully observable vector, bypassing the Clue Task. Both (a) and Mortar Mayhem Grid (b) utilize grid-like locomotion. (c) represents the default Mortar Mayhem task. Mystery Path Grid operates on grid-like locomotion, with (a) rendering the the goal and origin to the agent’s observation. (b) and (c) hide these, while (c) is the default Mystery Path task. Searing Spotlights (g) is not varied, while all conducted runs leverage observation reconstruction (Obs. Rec.) in this environment.
Before assessing GRU and TrXL on the default challenges of the finite environments, the first set of experiments focuses on confirming that the agents trained on Memory Gym indeed require memory. To accomplish this, we train several minor baselines alongside GRU and TrXL on simplified versions of the environments. These baselines include PPO without memory, PPO with frame stacking (4 and 16 grayscale frames), and PPO with absolute positional encoding (Vaswani et al., 2017) provided as vector observation.
Results on Mortar Mayhem (Figures 66(a)-6(c)) show the number of commands that were properly executed during one episode. In Mortar Mayhem Act Grid, where the commands are completely provided as vector observation, PPO without memory is ineffective, while the frame stacking agents achieve nearly either 5 or 8 commands with a slight upward trend. Agents using GRU or TrXL are the most effective, completing all 10 commands, while the one with absolute positional encoding follows closely, completing 9.5 commands. The positional encoding baseline adds the episode’s current time step to the agent’s observation allowing for temporal awareness within the sequential execution of commands. Once the complete task is present (Figure 6(b)), only GRU and TrXL prove to be effective and therefore we conclude that Mortar Mayhem demands memory.
We obtain a similar impression when training on Mystery Path Grid where the origin and goal are not perceivable by the agent (Figure 6(e)). In this case, TrXL outperforms GRU in terms of sample efficiency, while both methods are effective. If the origin and the goal are visible, a horizon of 16 frames is sufficient to train an effective policy as shown by the frame stacking agent (Figure 6(d)). Stacking 4 frames or leveraging positional encoding leads to an IQM success rate of 40% with a slight trend upward. The memory-less agent’s success rate goes up to about 11%.
No success is proven by the minor baselines in Searing Spotlights as seen in Figure 6(g). Only in this experiment, the minor baselines also leverage observation reconstruction. Utilizing the observation reconstruction loss significantly improves the sample efficiency of GRU and TrXL. The reason for this can be found in rare events: the agent, the coin, and the exit are seldomly visible. Hence, the visual encoder takes advantage from the auxiliary learning signal provided by the observation reconstruction. However, both GRU variants, with and without observation reconstruction, exhibit notably higher sample efficiency compared to their respective TrXL counterparts.
When facing the default task of Mortar Mayhem, a different result becomes apparent (Figure 6(c)). TrXL with observation reconstruction solves all 10 commands after training for about 100 million steps, while GRU converges to only about 6 commands. GRU’s performance drops to about 5 commands if the observation reconstruction loss is used. If TrXL trains without this auxiliary loss, its performance drops to about only one successful command.
@@
In the case of Mystery Path (Figure 6(f)), we opted not to train agents using observation reconstruction, as the always visible agent is the only entity that requires visual encoding. In this scenario, TrXL achieves the desired success rate in approximately 100 million steps, whereas GRU requires twice as many steps to achieve the same level of performance. If wall-time efficiency is compared, however, TrXL and GRU are roughly on par (Annex D).
In summary, TrXL proves to be more effective than GRU in Mortar Mayhem and exhibits higher sample efficiency in the case of Mystery Path. However, it is worth noting that GRU, both with and without observation reconstruction, outperforms its respective TrXL counterparts in Searing Spotlights.
@@ 5.2 Endless Environments: GRU is more effective than Transformer-XL
To our surprise, across all three endless environments, the recurrent agent consistently outperforms the transformer-based agent by a significant margin. This diverges from the observations made by Parisotto et al. (2020) and Lampinen et al. (2021), where an LSTM was found to be less effective. Notably, the most substantial performance gap between GRU and TrXL emerges in the results of Endless Mortar Mayhem (Figure 7(a)). Within this context, GRU attains an impressive IQM of 125 executed commands using roughly 300 million steps, whereas TrXL only attains an IQM of 18 commands. Furthermore, the incorporation of observation reconstruction exacerbates this performance difference, leading TrXL to a further deterioration, resulting in an IQM of 10. Furthermore, GRU exhibits consistently longer IQM episode lengths, with TrXL achieving an IQM of 288 steps and GRU reaching a peak IQM episode length of 3462 steps.
The outcomes depicted in Figure 7(b) illustrate the results obtained from the Endless Mystery Path environment. In this case, GRU remains more effective than TrXL, although the gap between their performances is narrower compared to Endless Mortar Mayhem. GRU achieves an IQM undiscounted return of 2.65, with TrXL achieving 2.02. Although neither memory approach converges within the allocated training budget of 819 million steps, GRU displays a more pronounced upward trend in performance. In terms of episode length, neither agent achieves lengths as seen in in Endless Mortar Mayhem. GRU attains an IQM episode length of 600 steps in Endless Mystery Path, while TrXL reaches up to 400 steps. This divergence can be attributed to the episode’s dynamics, where agents tend to fall off after significant progress, necessitating more steps to regain lost ground and accumulate further rewards.
@@ 5.3 Investigating Transformer-XL’s Surprisingly Low Effectiveness
Figure 8: Experiments on Endless Mortar Mayhem with varied TrXL configurations. The learning rate (LR) is adjusted to decay from $42.75e-$ 4 to $41.0e-$ 4 over 160 million steps, compared to the previous $51.0e-$ 5. We also incorporate observation reconstruction (Obs. Rec.) and observe the greatest improvement based on a ground truth estimation head (GT) as sanity check. Another test augments TrXL’s query with relative positional encoding (QPos). The final experiment combines the optimized learning rate, augmented query, and ground truth estimation.
@@ 5.3.1 Inadequate Networks Capacity
Our TrXL baseline comprises 2.8 million trainable parameters, while GRU consists of 4.05 million, prompting the question of whether TrXL’s model architecture lacks the necessary capacity. To investigate this, we conducted experiments varying the number of blocks (2, 3, 4), the embedding dimension size (256, 384, 512), and the memory window length (256, 384, 512). None of these adjustments led to closing the performance gap to GRU. Scaling up TrXL also increases its demand for GPU memory. When scaling up multiple architecture details, our implementation and hyperparameters may cause the training to exceed the available 40GB GPU memory of an NVIDIA A100. Workarounds include reducing the batch size or transferring training data between the GPU and CPU, but these options significantly worsen wall-time efficiency. Furthermore, another indication of untapped capacity lies in the amplification of the learning signal, which we elaborate on in the following section.
@@ 5.3.3 Lack of Temporal Information in the Initial Query
The next hypothesis pertains to the initial query of the first TrXL block. This query is constructed solely based on the features extracted from the observation encoders at the current time step, encompassing only minimal temporal information. In contrast, the query of the second block draws from the aggregated outcome of the memory window, thus capturing more substantial temporal information. However, we believe that the initial query could be further enriched with additional information. In Figure 8, one experiment augments the query with absolute positional encoding, providing direct access to the current time step. Despite this enhancement, the agent’s performance only moderately improves, reaching an IQM of 19 completed commands. We also explored embedding the original query with the previous output of the last TrXL block. However, this simple approach did not yield positive results.
@@ 5.3.4 Harmful Off-Policy Data in the Episodic Memory
Next, it can be discussed whether the combination of PPO and TrXL’s episodic memory pose a threat. PPO is an on-policy algorithm that consumes the training data for a few epochs. Once the second training epoch on the same batch commences, the batch is already considered off-policy. However, PPO’s loss function can mitigate this issue and hence allows for several epochs. But this mitigation is likely limited and we wonder whether this issue is enlarged by the episodic memory. In our current training setup, the agent collects 512 steps of data. If an episode exceeds 512 steps and is still running, the earlier steps from that episode become outdated over multiple epochs. In future work, it needs to be investigated if a stale episodic memory, which is based on off-policy data, hinders training progress. Frequently refreshing hidden states, a process used in the off-policy algorithm R2D2 that relies on recurrence (Kapturowski et al., 2019), is expected to be computationally inefficient.
@@ 5.4 Recurrence is not Vulnerable to Spotlight Perturbations
Several comprehensive studies have delved into such nuances of PPO’s implementation (Engstrom et al., 2020; Andrychowicz et al., 2021; Huang et al., 2022b). Among them, only Andrychowicz et al. (2021) turned off advantage normalization and found negligible performance impact. However, their extensive study focused on control tasks with undiscounted returns in the thousands, contrasting Memory Gym’s finite environments where the undiscounted returns do not exceed 2. During hyperparameter optimization, turning off advantage normalization also enhanced performance in Mortar Mayhem and Mystery Path. Given the critical importance of utilizing raw advantages, there remains an opportunity to delve even deeper into PPO to comprehend and refine its essential aspects.

### v5

@@ 5 Experimental Analysis
We run empirical experiments on Memory Gym using the just introduced TrXL baseline and one based on GRU, both powered by Proximal Policy Optimization (Schulman et al., 2017). The next subsection explains why GRU and TrXL are the only baselines considered. Next, we detail the evaluation protocol, drawing inspiration from Agarwal et al. (2021). Section 5.3 presents results on the finite tasks, showing that TrXL exhibits stronger policies than GRU in Mortar Mayhem, while being more sample efficient in Mystery Path. Searing Spotlights is an exception, where TrXL is less sample efficient. Contradicting our initial expectations and the results on the finite environments, we unveil surprising findings that highlight GRU’s superiority over TrXL in all endless environments. Afterward, we compare the results from the finite and endless environments, demonstrating that the finite settings offer only a limited view of the agents’ true memory capabilities. We then broadly explore several hypotheses on why TrXL might be less effective. Finally, we falsify that recurrent agents are vulnerable to spotlight perturbations when using enhanced hyperparameters (e.g., not normalizing advantage estimates).
5.1 Considered Baselines
@@
Other related and significant baselines, such as the Memory Recall Agent (MRA) (Fortunato et al., 2019), Hierarchical Chunk Attention Mechanism (HCAM) (Lampinen et al., 2021), and Gated Transformer-XL (GTrXL) (Parisotto et al., 2020), are not considered due to their closed-source nature, which hampers reproduction. Unlike our previous study (Pleines et al., 2023), we exclude HELM (History comprEssion via Language Models) (Paischer et al., 2022a) and its successor, HELMv2 (Paischer et al., 2022b). This decision stems from HELM’s suboptimal wall-time efficiency, which is six times more expensive than that of a GRU agent (Pleines et al., 2023), making hyperparameter optimization impractical for our purposes. Given our strict compute budget of 25,000 GPU hours, we dedicated our resources to developing the endless environments and the TrXL baseline instead.
For a comprehensive overview of the hyperparameters used in our experiments, see Appendix B. We also conduct preliminary tests using GTrXL and Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) on the finite environments without extensive hyperparameter tuning. GTrXL is originally based on the DRL algorithm V-MPO (Song et al., 2020). We modify our TrXL baseline to adopt the gating mechanism. Results for GTrXL and LSTM on the finite environments are provided in Appendix D. These models usually underperformed when applied directly, suggesting that out-of-the-box applications might necessitate hyperparameter optimization.
Additionally, we reaffirm that Memory Gym’s environments require memory by training naive baselines, such as frame stacking (Appendix C). A selection of wall-time efficiency statistics is available in Appendix E.
5.2 Evaluation Protocol
@@
A major challenge in DRL research is the statistical uncertainty resulting from high computational costs and the stochastic nature of training algorithms. Due to these expenses, many studies rely on limited data, such as those evaluating PPO, MRA, and HCAM with only three runs per experiment. Under these conditions, Agarwal et al. (2021) propose more robust statistical tools to reliably evaluate the performance of agents trained with DRL. By following their guidelines, we aggregate 5 training runs using the interquartile mean (IQM), which is preferred because the mean is sensitive to outliers and the median requires much more data. To measure and visualize variability, the following sample efficiency curves depict the stratified bootstrapped confidence interval of 0.95 (represented by shaded areas on the plots). To evaluate the generalization ability of our baselines, we extensively test them on 50 novel environment seeds that were excluded from the training process. This evaluation is repeated 3 times for each data point, resulting in a comprehensive assessment based on a total of 750 episodes for every data point presented.
@@ 5.3 Finite Environments: Transformer-XL is often more efficient than GRU
Figure 6: Performance comparison across instances of Memory Gym’s finite environments. Notably, GRU and TrXL are primarily different in terms of sample efficiency. Only in Mortar Mayhem (c), levels of effectiveness are apparent. Mortar Mayhem Act Grid (a) provides commands as a fully observable vector, bypassing the Clue Task. Both (a) and Mortar Mayhem Grid (b) utilize grid-like locomotion. (c) represents the default Mortar Mayhem task. Mystery Path Grid operates on grid-like locomotion, with (d) rendering the goal and origin to the agent’s observation. (e) and (f) hide these, while (f) is the default Mystery Path task. Searing Spotlights (g) is not varied, while the baselines may leverage observation reconstruction (Obs. Rec.) in this environment.
The introduction of the TrXL baseline motivates conducting benchmarks on the finite environments of Memory Gym, as depicted in Figure 6. Due to the upper bound of the finite tasks, the differences between GRU and TrXL are primarily evident in terms of sample efficiency, with Mortar Mayhem 66(c) being the only exception. We will address this specific scenario after explroing efficiency trends.
Results on Mortar Mayhem (Figures 66(a)-6(c)) show the number of commands that were properly executed during one episode. In Mortar Mayhem Act Grid, GRU and TrXL complete all 10 commands, while TrXL is slightly less efficient. Once the clue task is present (Figure 6(b)), TrXL is more efficient than GRU.
@@
In the Mystery Path environments, we opted not to train agents using observation reconstruction, as the always visible agent is the only entity that requires visual encoding. GRU is only more sample efficient in Mystery Path Grid where the origin and goal are perceivable by the agent (Figure 6(d)). In the other two scenarios (Figures 66(e)-6(f)), TrXL needs less samples to achieve the maximum success rate. If wall-time efficiency is compared, TrXL and GRU are roughly on par (Annex E).
In Searing Spotlights (Figure 6(g)), utilizing the observation reconstruction loss is key. It significantly improves the sample efficiency of GRU and TrXL. The reason for this can be found in rare events: the agent, the coin, and the exit are seldomly visible. Hence, the visual encoder takes advantage from the auxiliary learning signal provided by the observation reconstruction. Nevertheless, both GRU variants, with and without observation reconstruction, exhibit notably higher sample efficiency compared to their respective TrXL counterparts.
Returning to the exception noted in Mortar Mayhem (Figure 66(c)), we quantify effectiveness levels only for GRU, GRU with observation reconstruction, and TrXL. GRU completes 6 commands, observation reconstruction reduces GRU’s performance to 5, while TrXL without this auxiliary head performs poorly, completing only 1 command. Conversely, TrXL with observation reconstruction effectively resolves all 10 commands, though its true limit remains uncertain. As previously argued in Section 3.1, increasing the difficulty is necessary to accurately assess TrXL’s limits, which we will further investigate by exploiting the automatic curriculum of the endless environments.
@@ 5.4 Endless Environments: GRU is more effective than Transformer-XL
To our surprise, across all three endless environments, the recurrent agent consistently proves to be much more effective than the transformer-based agent. This diverges from the observations made by Parisotto et al. (2020) and Lampinen et al. (2021), where an LSTM was found to be less effective. This subsection presents the raw results, and the following one discusses how finite tasks provide only a truncated view of the memory capabilities of the evaluated agents.
Notably, the most substantial gap between GRU and TrXL emerges in the results of Endless Mortar Mayhem (Figure 7(a)). Within this context, GRU attains an impressive IQM of 125 executed commands using roughly 300 million steps, whereas TrXL only attains an IQM of 18 commands. The incorporation of observation reconstruction exacerbates this performance difference, leading TrXL to a further deterioration, resulting in an IQM of 10. GRU exhibits consistently longer IQM episode lengths, with TrXL achieving an IQM of 288 steps and GRU reaching a peak IQM episode length of 3462 steps.
The outcomes depicted in Figure 7(b) illustrate the results obtained from the Endless Mystery Path environment. In this case, GRU remains more effective than TrXL, although the gap between their performances is narrower compared to Endless Mortar Mayhem. GRU achieves an IQM undiscounted return of 2.65, with TrXL achieving 2.02. Although neither memory approach converges within the allocated training budget of 819 million steps, GRU displays a more pronounced upward trend in performance. In terms of episode length, neither agent achieves lengths as seen in Endless Mortar Mayhem. GRU attains an IQM episode length of 600 steps in Endless Mystery Path, while TrXL reaches up to 400 steps. The divergence between the returns and the episode lengths are attributed to the episode’s dynamics, where agents tend to fall off after significant progress, necessitating more steps to regain lost ground and accumulate further rewards.
@@
5.5 Memory Effectiveness: Finite Boundaries vs. Endless Potential
In this subsection, we highlight that the insights derived from finite environments provide only a truncated view of the true memory capabilities of the agents benchmarked. To recall, higher memory effectiveness implies the ability to memorize more information over extended periods.
Mortar Mayhem stands as an exception among finite environments, where varied levels of effectiveness are observable. One might assume that these differences would align with the results from endless tasks, yet this alignment does not occur. Notably, in the endless setting, GRU completes significantly more commands than those 6 observed in the finite environment. This increased quantity can be attributed to two key factors. Firstly, the environment’s curriculum begins at an easier level by only demanding one command initially. Secondly, Endless Mortar Mayhem allows for two strategies for memorizing commands: the direct visualization of the command and the agent’s ability to infer commands based on its past actions. This may allow for the agent’s policy to adapt faster to growing difficulties.
In the Mystery Path environment, challenges are limited to a maximum path length of 17 tiles, providing only a basic indication of the memory requirements. This measure is approximate, as the procedurally generated path may include multiple turns, increasing the likelihood of the agent falling off. Both GRU and TrXL effectively master this challenge, but they actually surpass these capabilities in the endless version. The undiscounted return reflects the length of the path to the farthest point reached by the agent: 26.5 tiles for GRU and 20 tiles for TrXL, indicating that both models are capable of retaining and recalling more information than demonstrated in the finite setting. As scores correlate with episode duration, agents must maintain robust memory over prolonged periods.
Quantifying performance in Searing Spotlights involves navigating to two entities (coin and exit), while the agent must maintain its hidden position. In finite tasks, effective episodes typically last fewer than 50 steps. However, in Endless Searing Spotlights, GRU and TrXL can collect 20 and 6 coins and maintain their hidden positions for 600 and 200 steps, respectively. This performance demonstrates that both agents not only achieve higher scores but also sustain their concealed positions for longer durations than observed in the finite version of Searing Spotlights.
To conclude, endless tasks reveal more effective memory capabilities in both baselines beyond what is evident in finite tasks.
@@ 5.6 Investigating Transformer-XL’s Surprisingly Low Effectiveness
Figure 8: Experiments on Endless Mortar Mayhem with varied TrXL configurations. The learning rate (LR) is adjusted to decay from $2.75$e-$4$ to $1.0$e-$4$ over 160 million steps, compared to the previous $1.0$e-$5$. We also incorporate observation reconstruction (Obs. Rec.) and observe the greatest improvement based on a ground truth estimation head (GT) as sanity check. Another test augments TrXL’s query with absolute positional encoding (QPos). The final experiment combines the optimized learning rate, augmented query, and ground truth estimation.
@@ 5.6.1 Inadequate Networks Capacity
Our TrXL baseline comprises 2.8 million trainable parameters, while GRU consists of 4.05 million, prompting the question of whether TrXL’s model architecture lacks the necessary capacity. To investigate this, we conducted experiments varying the number of layers (2, 3, 4), the embedding dimension size (256, 384, 512), and the memory window length (256, 384, 512). None of these adjustments led to closing the performance gap to GRU. Even when the GRU cell is scaled down to a dimension of 384, totaling 2.7 million parameters, it still surpasses TrXL by a large margin.
Scaling up TrXL also increases its demand for GPU memory. When scaling up multiple architecture details, our implementation and hyperparameters may cause the training to exceed the available 40GB GPU memory of an NVIDIA A100. Workarounds include reducing the batch size or transferring training data between the GPU and CPU, but these options significantly worsen wall-time efficiency. Furthermore, another indication of untapped capacity lies in the amplification of the learning signal, which we elaborate on in the following section.
@@ 5.6.3 Lack of Temporal Information in the Initial Query
The next hypothesis pertains to the initial query of the first TrXL layer. This query is constructed solely based on the features extracted from the observation encoders at the current time step, encompassing only minimal temporal information. In contrast, the query of the second layer draws from the aggregated outcome of the memory window, thus capturing more substantial temporal information. However, we believe that the initial query could be further enriched with additional information. In Figure 8, one experiment augments the query with absolute positional encoding, providing direct access to the current time step. Despite this enhancement, the agent’s performance only moderately improves, reaching an IQM of 19 completed commands. We also explored embedding the original query with the previous output of the TrXL block. However, this simple approach did not yield positive results.
@@ 5.6.4 Harmful Off-Policy Data in the Episodic Memory
Next, it can be discussed whether the combination of PPO and the cached hidden states of TrXL’s episodic memory pose a threat. PPO is an on-policy algorithm that consumes the training data for a few epochs. Once the second training epoch on the same batch commences, the batch is already considered off-policy. However, PPO’s loss function can mitigate this issue and hence allows for several epochs. But this mitigation is likely limited and we wonder whether this issue is enlarged by the cached hidden states. In our current training setup, the agent collects 512 steps of data. If an episode exceeds 512 steps and is still running, the older hidden states from that episode become outdated over multiple epochs. To address this, future research should investigate the impact of stale hidden states on training efficacy. One potential strategy could involve computing more fresh hidden states alongside utilizing cached ones. Alternatively, cached hidden states could be periodically recomputed, albeit at the expense of increased inference overhead. Drawing parallels, Kapturowski et al. (2019) explored similar challenges with their recurrent and off-policy algorithm R2D2.
@@ 5.7 Recurrence is not Vulnerable to Spotlight Perturbations
Several comprehensive studies have delved into such nuances of PPO’s implementation (Engstrom et al., 2020; Andrychowicz et al., 2021; Huang et al., 2022). Among them, only Andrychowicz et al. (2021) turned off advantage normalization and found negligible performance impact. However, their extensive study focused on control tasks with undiscounted returns in the thousands, contrasting Memory Gym’s finite environments where the undiscounted returns do not exceed 2. During hyperparameter optimization, turning off advantage normalization also enhanced performance in Mortar Mayhem and Mystery Path. Given the critical importance of utilizing raw advantages, there remains an opportunity to delve even deeper into PPO to comprehend and refine its essential aspects.